ltvers.dtx

Johannes Braams  David Carlisle  Alan Jeffrey
Leslie Lamport  Frank Mittelbach  Chris Rowley

Rainer Schöpf

1995/05/12

1 Version Identifictation

First we identify the date and version number of this release of L^{A}TEX, and set \everyjob so that it is printed at the start of every L^{A}TEX run.

\fmtname \fmtversion

1 ⟨∗2ekernel⟩

 2 \def\fmtname{LaTeX2e}
 3 \edef\fmtversion{1995/12/01}

Check that the format being made is not too old.

 4 \def\reserved@a#1/#2/#3\@nil{%
 5 \count@\year
 6 \advance\count@-#1\relax
 7 \multiply\count@ by 12\relax
 8 \advance\count@\month
 9 \advance\count@-#2\relax}
10 \expandafter\reserved@a\fmtversion\@nil

\count@ is now the age of this file in months.

11 \ifnum\count@>12
12 \typeout{^^J%
13 ******************************************************************^^J%
14 *  You are attempting to make a LaTeX format from a source file^^J%
15 *  That is more than one year old.^^J%
16 *^^J%
17 *  If you enter <return> to scroll past this message then the format^^J%
18 *  will be built, but please consider obtaining newer source files^^J%
19 *  before continuing to build LaTeX.^^J%
20 *^^J%
21 *  LaTeX is re-issued every 6 months June and December.^^J%
22 ******************************************************************^^J%

23 }

24 \errhelp{To avoid this error message, obtain new LaTeX sources.}
25 \errmessage{LaTeX source files more than 1 year old!}

26 \fi
27 \let\reserved@a\relax

This startup banner may be further modified by the code in ltfinal.dtx if a patch file is present.

28 \everyjob{\typeout{\fmtname
29 ⟨autoload⟩\space(autoload version)%
30  \space<\fmtversion>}}
31 \immediate\write16{\fmtname
32 ⟨autoload⟩\space(autoload version)%
33  \space<\fmtversion>}

34 ⟨/2ekernel⟩